A board measuring 120 inches long needs to be cut into two pieces so that one piece will be four times longer than the other. Find the lengths of the two pieces.
This is what I done so far I found the length of one piece.
x+4=120
5x=120
x=24
Now I just need to find out how to solve for the other length

Found 2 solutions by Earlsdon, TakeATuition.com:
Answer by Earlsdon(6294)

Congratulations! You have correctly identified the length of one piece.Please note however that your first equation should have been:
x+4x = 120
5x = 120
x = 24
Now to find the other length, you can either subtract the 24 from the original length of 120 to get 120-24 = 96 or, since you know that one piece is x and the other piece is 4x and x = 24, then 4x = 4(24) = 96.

Answer by TakeATuition.com(57)

SOLUTION -
Let the length of one board be x inches
Therefore the length of other board is 4x (since one piece is 4 times the other)
now sum of both pieces = 120
x + 4x = 120
5x = 120
5x/5 = 120/5

x = 24
now
one piece ix x inches = 24 inches
another piece is 4x = 4*24 = 96 inches
FOR MORE HELP VISIT : www.TakeATuition.com